# Flowr.root – A flow matching based foundation model for joint multi-purpose structure-aware 3D ligand generation and affinity prediction

### v3

@@ 4 Methods
Flowr.root is built on the recently proposed Flowr architecture, an -equivariant flow matching model that learns a mixed continuous/discrete transport map. This map transforms a prior distribution (e.g., random noise or fragment anchors for coordinates) to the target ligand distribution within a given protein pocket (model:flowr). The model consists of two main components: a pocket encoder and a ligand decoder.
@@
Protein pockets are extracted from holo structures by cutting residues within a defined radius around the binding site. These pockets are then encoded using an equivariant self-attention module followed by equivariant feed-forward layers, generating a set of invariant and equivariant protein features (model:flowr). The ligand decoder processes noisy ligand coordinates, partial charges, atom types, bond orders, and hybridization states through equivariant self-attention modules, which capture intra-ligand dependencies. A cross-attention layer then integrates contextual information from the pocket features to the hidden ligand features. For additional details, we refer to model:flowr.
In Flowr.root, the ligand decoder has three output heads: (1) a structure head that predicts atomic coordinates, atom types, bond orders, charges, and hybridization states; (2) an affinity head that estimates four binding affinity types—IC50, pKi, pKd, and pEC50; and (3) a confidence head that provides uncertainty estimation based on pLDDT (folding:alphafold3).
This multi-task formulation ensures that structural accuracy, potency estimation, and confidence calibration occur simultaneously during ligand generation. Following model:flowr, Flowr.root also supports three generative modes within a single backbone: (1) de novo generation, (2) interaction-guided generation (to preserve or enforce pharmacophoric patterns), and (3) fragment-based elaboration. The latter mode uses fragments as anchors to generate new structures that are consistent with the binding pocket, enabling fragment growing, merging, linking, and scaffold hopping/elaboration. This multi-mode capability allows flexible applications in hit identification to fragment-based lead optimization.
@@
The core challenge in protein-ligand modeling, however, lies not just in generation, but in accurately predicting protein-ligand interactions and potency. The sparsity of experimental binding data—where most ligand–protein combinations remain unexplored—poses a significant obstacle to improving these predictions. While broad-coverage datasets encompass a wide spectrum of protein targets, they are intrinsically sparse, with few ligands tested systematically across different biological contexts. This imbalance leads to biased datasets, which explains why 3D machine and deep learning models—often relying only on protein or ligand descriptors and ignoring interaction patterns—can perform surprisingly well on affinity prediction benchmarks like PDBBind. These inflated results also reveal systematic biases in the data, with activity clusters within certain protein families or chemical series (volkov2022frustration).
Real-world drug discovery requires understanding nuanced structure–activity relationships (SAR) across hundreds or even thousands of related molecules within the same binding site—a richness uniquely captured by dense bioactivity landscapes. These landscapes encode redundancy and critical phenomena like activity cliffs, where small structural changes can lead to significant shifts in bioactivity due to interaction changes.
This motivates the multi-stage training approach of Flowr.root: starting with sparse, lower-fidelity data to learn broad 3D chemistry and binding principles, and progressively refining the model with denser, higher-fidelity and project-specific datasets that capture subtle SAR patterns essential for effective hit expansion to lead optimization. This multi-stage pipeline is intended to closely mirrors the practical needs of early drug discovery campaigns, addressing the limitations imposed by data sparsity in current 3D affinity prediction models.
In summary, the Flowr.root framework consists of three stages:
@@
Protein–ligand complexes with and without affinity labels (2.5M datapoints from Plinder, BindingMoad, BindingNet, SAIR, KIBA-3D, Kinodata-3D, OMol25), which capture protein-ligand structural diversity, in parts annotated with affinity data. This dataset is split into lower- and higher-fidelity categories. We classify complexes with computational origin (i.e., BindingNet, SAIR, KIBA-3D, Kinodata-3D) as lower-fidelity and those stemming from co-crystal data (i.e., Plinder, BindingMoad) as higher-fidelity, with the exception of the high-confidence subset of BindingNet, which is treated as higher-fidelity.
Stage 2: High-fidelity fine-tuning.
Pre-trained weights are adapted to thoroughly curated, drug-like, high-quality affinity datasets. In this work, we use SPINDR and HiQBind (30k complexes) as datasets for multi-target fine-tuning. However, the main goal of this stage is to fine-tune Flowr.root on high-quality in-house data, which teaches the model the fine details of protein-ligand interactions and binding affinity.
@@
This ensures that Flowr.root can bridge the gap between large-scale, noisy public data and small-scale, high-quality, in-assay data, producing distributions that align with specific discovery campaigns. Different domain adaptation strategies are used to fine-tune and steer the pre-trained Flowr.root model:
@@
Direct preference alignment: The model is fine-tuned with project-specific preferences, such as avoiding known liability motifs or penalizing activity cliffs (future work).
Standard fine-tuning: Parameter-efficient fine-tuning, with the freezing of non-critical modules, helps avoid catastrophic forgetting while aligning the model to project-specific structure-activity modes.
4.2 Training and Inference
Closely following the methodology of model:flowr, we adopted the training and parameterization scheme from FLOWR.multi. Specifically, we employed a 4-layer pocket encoder with and a 12-layer ligand decoder with . The equivariant feature dimension was set to for both the pocket encoder and ligand decoder. For latent attention, we utilized a latent size of 64 with 32 attention heads.
To further enhance model stability and generalization, we incorporated residual coordinate and subsequent edge updates using Gaussian radial basis expansions (model:eqgatdiff), applied to the layer-wise latent message-passing outputs of the Flowr model backbone (model:flowr). Additionally, we included distance and cross-product computations between latent equivariant node features within the message-passing layers, which we found to improve model performance. Specifically, in the pairwise interactions between nodes, given the equivariant queries and keys , we define the pairwise difference tensor as
@@ 4.2 Training and Inference
The model structure output comprises predictions for atomic coordinates, atom types, charge, hybridization, and bond types. The overall loss function is:
with 33.8M trainable parameters.
@@
Flowr.root jointly models both continuous (coordinates) and discrete (atom types, bond orders, charges, hybridizations) molecular features. For coordinates, we employ continuous flow matching (cfm:minibatch-ot), while discrete flow models (model:multiflow) are used for categorical variables. The model is trained to recover the original ligand from a noisy ligand by learning the conditional distribution , minimizing mean squared error for coordinates and cross-entropy loss for categorical features. Given a pocket structure , novel ligands are generated by iteratively refining an initial noisy ligand . The generative process follows a learned vector field for continuous features and a discrete integration scheme for categorical attributes (model:multiflow; model:flowr).
Binding affinity prediction
@@
In drug discovery, a variety of parameters, including Ki, Kd, IC50, and EC50, are reported as potency measures. These parameters differ in their definitions and experimental setups, which complicates their direct comparison. For example, Kd is an equilibrium constant that directly measures the strength with which a ligand binds to its target protein. On the other hand, a Ki is an equilibrium constant indicating how well a given inhibitor inhibits the binding of a natural substrate. In other words, while Kd measures the ”stickiness” of a molecule to a target, Ki measures how well the inhibitor blocks the enzyme from its natural substrate. Though these values often correlate, they are not always interchangeable. Likewise, EC50 and IC50 are used for biological assays. Though they correlate with Kd and Ki, their interpretation depends on assay conditions.
@@
Therefore, we predict each potency measure separately, aligning better with the complexity of drug-target data. Given invariant and equivariant ligand features , pocket features , and ligand-pocket interaction features extracted from the ligand decoder’s latent message-passing module, the binding affinity head computes:
@@
where is a Gated equivariant block (model:painn) to combine equivariant and invariant feature tensors, and is a multi-layer perceptron with two linear layers and SiLU activation. The input to is the concatenation of normalized invariant features, where , and denotes the number of atoms in the ligand. The concatenated feature tensor is passed to task-specific heads:
@@
The affinity prediction loss is computed using the mean Huber loss between predicted and true affinity values in log units (pIC50, pKd, pKi, pEC50):
@@
Using distance thresholds and L1 distance errors , the LDDT score for each ligand atom is:
@@
4.3 Inference-time scaling
Inference-time scaling is a powerful approach that allocates additional computation during generation, rather than solely relying on larger architectures (snell2024scalingllmtesttimecompute; setlur2024rewardingprogressscalingautomated). Originally developed for large language models, this approach has been extended to diffusion and flow matching models. Diffusion models leverage stochasticity through particle sampling (wu2023practical; model:pilot) and Feynman-Kac steering (singhal2025a), where multiple trajectories are generated and resampled based on reward functions. Flow matching models, despite being deterministic, achieve inference-time scaling through SDE-based generation, interpolant conversion, and Rollover Budget Forcing (RBF) for adaptive resource allocation (kim2025inferencetimescalingflowmodels), improving upon diffusion-based advances (ma2025inferencetimescalingdiffusionmodels).
@@ 4.3 Inference-time scaling
The sequential Monte Carlo (SMC) method (smc_del_moral_doucet) extends this concept to a time-sequential setting by maintaining particles and updating their importance weights over time (wu2023practical). A straightforward weighting scheme involves evaluating the current batch of particles at time , computing their scores , and obtaining resampling weights through softmax normalization: . Particles are then resampled from the discrete distribution with probability proportional to their weights, as done in PILOT (model:pilot).
4.4 Quantum Mechanical Calculations on Protein-Ligand Complexes
@@
To benchmark Flowr.root’s affinity prediction head, we perform quantum mechanical (QM) calculations on protein-ligand complexes and correlate the results with predicted affinities. Calculations were conducted using ULYSSES (Menezes:ULYSSES), a general-purpose semi-empirical library offering multiple molecular Hamiltonians. We employ GFN2-xTB (xtb_gfn2) with ALPB aqueous solvation (alpb). GFN2-xTB is a simplified density functional method that parametrizes interparticle interaction integrals against reference data, significantly accelerating calculations (xtb_gfn2)—a critical advantage for large atomic systems such as protein-ligand complexes. Additionally, GFN2-xTB provides a balanced treatment of electrostatic interactions and incorporates state-of-the-art dispersion corrections, enabling accurate descriptions of non-bonded complexes.

### v5

@@ Methods
Flowr.root is built on the recently proposed Flowr architecture, an -equivariant flow matching model that learns a mixed continuous/discrete transport map. This map transforms a prior distribution (for example, random noise or fragment anchors for coordinates) to the target ligand distribution within a given protein pocket [17]. The model consists of two main components: a pocket encoder and a ligand decoder.
Protein pockets are extracted from holo structures by cutting residues within a defined radius around the binding site. We use a 7Å cutoff radius, which represents a practical compromise between computational tractability and interaction coverage. This cutoff fully encompasses hydrogen bonds directly relevant to ligand binding (which decay rapidly beyond 5.5–6Å), captures vibrational entropy contributions from direct protein-ligand interactions, and includes conformational entropy changes in the binding pocket. While long-range electrostatic effects are system-dependent (primarily affecting charged ligands), the 7Å cutoff provides sufficient coverage for neutral ligands and partial cancellation occurs for charged species. Water-mediated interactions are treated implicitly, consistent with many previous ML approaches. The cutoff is a configurable parameter that can be adjusted at both training and inference time.
@@
These pockets are then encoded using an equivariant self-attention module followed by equivariant feed-forward layers, generating a set of invariant and equivariant protein features [17]. The ligand decoder processes noisy ligand coordinates, partial charges, atom types, bond orders, and hybridization states through equivariant self-attention modules, which capture intra-ligand dependencies. A cross-attention layer then integrates contextual information from the pocket features to the hidden ligand features. The pocket encoder processes features on a full-atom basis through -equivariant self-attention layers using invariant residue embeddings, equivariant coordinate features, and pairwise fully-connected distance information between all pocket atoms. Via fully-connected, equivariant cross-attention, the ligand decoder interacts with the invariant and equivariant pocket encodings for ligand generation. For additional details, we refer to Cremer et al. [17].
In Flowr.root, the ligand decoder has three output heads: (1) a structure head that predicts atomic coordinates, atom types, bond orders, charges, and hybridization states; (2) an affinity head comprising four separate potency and binding affinity heads—specifically, for IC50, pKi, pKd, and pEC50 prediction—and (3) a confidence head that provides uncertainty estimation based on pLDDT [44]. Importantly, these heads serve distinct yet complementary purposes: the structure head generates geometrically valid ligand conformations within the protein pocket, while the affinity head provides potency and binding affinity estimates that enable (1) in-distribution evaluation of generated structures, (2) inference-time steering toward higher-affinity compounds via importance sampling, and (3) domain adaptation of both structure and affinity prediction through joint finetuning. This joint training paradigm thus facilitates coherent, in-distribution property-guided generation without requiring external scoring functions—avoiding distribution mismatch and additional computational overhead—while supporting effective domain adaptation.
Importantly, Flowr.root builds upon the multi-task training framework of Flowr [17], yet extends it substantially to support four generative modes within a single backbone: (1) de novo and, (2) interaction-conditional generation for preserving or enforcing pharmacophoric patterns, (3) scaffold hopping and elaboration, and (4) fragment growing and replacement. Notably, with the latter two Flowr.root supports specifying which atoms to replace while preserving the remainder of the molecule. To enforce locality in these partial modifications, we introduce a flexible prior placement strategy that shifts the generative prior from the zero center-of-mass to the local replacement site. This multi-mode capability enables flexible application across hit identification and hit-to-lead and lead optimization with targeted structural modifications.
@@
The core challenge in protein-ligand modeling, however, lies not just in generation, but in accurately predicting protein-ligand interactions and potency. The sparsity of experimental binding data—where most ligand–protein combinations remain unexplored—poses a substantial obstacle to improving these predictions. While broad-coverage datasets encompass a wide spectrum of protein targets, they are intrinsically sparse, with few ligands tested systematically across different biological contexts. This imbalance leads to biased datasets, which explains why 3D machine and deep learning models—often relying only on protein or ligand descriptors and ignoring interaction patterns—can perform surprisingly well on affinity prediction benchmarks like PDBbind [111]. Real-world drug discovery requires understanding nuanced SAR across hundreds or even thousands of related molecules within the same binding site—a richness uniquely captured by dense bioactivity landscapes. These landscapes encode redundancy and critical phenomena like activity cliffs, where small structural changes can lead to substantial shifts in bioactivity due to interaction changes. This motivates the multi-stage training approach of Flowr.root: starting with sparse, lower-fidelity data to learn broad 3D chemistry and binding principles, and progressively refining the model with denser, higher-fidelity and project-specific datasets that capture subtle SAR patterns essential for effective hit expansion to lead optimization. This multi-stage pipeline is intended to closely mirror the practical needs of early drug discovery campaigns, addressing the limitations imposed by data sparsity in current 3D affinity prediction models. The Flowr.root framework consists of three stages:
We position Flowr.root as a foundation model in the sense formalized by Bommasani et al. [63], who define a foundation model as a model that is trained on broad data at scale meant to be adapted (for example, LoRA-fine-tuned) to a range of downstream tasks. Critically, foundation models are described as unfinished, intermediary assets requiring adaptation [63, 64, 65]. In this work, the foundation model property of Flowr.root refers specifically to its broad multi-stage pre-training on diverse chemical and structural data at scale, combined with its adaptability to specific downstream drug discovery tasks via parameter-efficient fine-tuning (LoRA; [47]). We note that zero-shot generalization especially across arbitrary bioactivity landscapes is not implied; rather, we expect the model to provide transferable representations that enable rapid, data-efficient domain adaptation.
@@
Protein–ligand complexes with and without affinity labels (2.5M datapoints from Plinder, BindingMoad, BindingNet, SAIR, KIBA-3D, Kinodata-3D, OMol25), which capture protein-ligand structural diversity, in parts annotated with affinity data. This dataset is split into lower- and higher-fidelity categories. We classify complexes with computational origin (that is, BindingNet, SAIR, KIBA-3D, Kinodata-3D) as lower-fidelity and those stemming from co-crystal data (that is, Plinder, BindingMoad) as higher-fidelity, with the exception of the high-confidence subset of BindingNet, which is treated as higher-fidelity.
Stage 2: High-fidelity finetuning.
Pre-trained weights are adapted to thoroughly curated, drug-like, high-quality affinity datasets. In this work, we use Spindr and HiQBind (30k complexes) as datasets for multi-target finetuning. However, this stage can be used to finetune Flowr.root on diverse, high-quality in-house data, if available, which teaches the model the fine details of protein-ligand interactions and binding affinity.
@@
This ensures that Flowr.root can bridge the gap between large-scale, noisy public data and small-scale, high-quality, dense in-assay data, producing distributions that align with specific discovery campaigns. Different domain adaptation strategies are used to finetune and steer the pre-trained Flowr.root model:
@@
Direct preference alignment: The model is finetuned with project-specific preferences, such as avoiding known liability motifs or penalizing activity cliffs (future work).
Finetuning: Parameter-efficient finetuning, namely Low-Rank Adapation (LoRA) [47], helps avoid catastrophic forgetting while aligning the model to project-specific structure-activity modes.
Training and Inference
Following the methodology of Cremer et al. [17], we adopted the training and parameterization scheme from FLOWR.multi. Specifically, we employed a 4-layer pocket encoder with and a 12-layer ligand decoder with . The equivariant feature dimension was set to for both the pocket encoder and ligand decoder. For latent attention, we utilized a latent size of 64 with 32 attention heads. Overall, Flowr.root has approximately 33 million trainable parameters. Throughout this work, the model has been trained on one Nvidia H100 node comprising eight H100 GPUs.
To further enhance model stability and generalization, we incorporated residual coordinate and subsequent edge updates using Gaussian radial basis expansions [39], applied to the layer-wise latent message-passing outputs of the Flowr model backbone [17]. Additionally, we included distance and cross-product computations between latent equivariant node features within the message-passing layers, which we found to improve model performance. Specifically, in the pairwise interactions between nodes, given the equivariant queries and keys , we define the pairwise difference tensor as
@@
The model structure output comprises predictions for atomic coordinates, atom types, charge, hybridization, bond types, bond lengths and bond angles. The overall loss function reads:
with denoting the respective loss weighting. To encourage geometrically accurate bond lengths, we compute the mean squared error between predicted and ground truth distances for all bonded atom pairs:
where denotes the set of bonded atom pairs extracted from the bond adjacency matrix, considering only the upper triangle to avoid redundancy. To preserve local molecular geometry, we penalize deviations in bond angles for all valid triplets where atom is bonded to both atoms and . For each triplet, the bond angle is computed as:
The bond angle loss is then defined using the Huber loss for robustness:
where is the set of valid angle triplets, and the Huber loss with threshold is defined as:
We use radians to balance sensitivity to small deviations while remaining robust to outliers.
@@
Flowr.root jointly models both continuous (coordinates) and discrete (atom types, bond orders, charges, hybridizations) molecular features. For coordinates, we employ continuous flow matching [66], while discrete flow models [42] are used for categorical variables. The model is trained to recover the original ligand from a noisy ligand by learning the conditional distribution , minimizing mean squared error for coordinates and cross-entropy loss for categorical features. Given a pocket structure , new ligands are generated by iteratively refining an initial noisy ligand . The generative process follows a learned vector field for continuous features and a discrete integration scheme for categorical attributes [42, 17].
Anisotropic and reference-ligand-based prior strategies
Standard flow matching employs an isotropic Gaussian as the source distribution . In partial generation modes—scaffold hopping, scaffold elaboration, linker inpainting, core growing, and fragment growing—a fixed molecular substructure provides geometric context that constrains where new atoms should be placed. An isotropic prior distributes probability mass without directional bias, forcing the learned velocity field to perform long-range, direction-dependent transport to reach the target region. A distinguishing property of flow matching, in contrast to diffusion models where the terminal distribution is determined by the noise schedule, is that need not be a standard Gaussian: any distribution from which efficient sampling is possible suffices [66]. The velocity field is trained to transport whatever is supplied toward , enabling per-sample prior selection and even mixed isotropic and anisotropic batches within a single training step. Flowr.root exploits this flexibility by constructing task-specific anisotropic Gaussian priors whose covariance encodes the spatial extent and directionality of the generation context.
Two complementary covariance constructions are employed. The shape-based covariance captures the spatial extent of a set of reference atom coordinates. Given a set of reference atom coordinates (variable atoms or the full molecule, depending on the generation mode), the centered coordinate matrix is
and the sample covariance is
Eigendecomposing with , the eigenvalues are first normalized to satisfy :
then clamped to and re-normalized to restore the trace constraint. The shape-based covariance is
This construction is applied in scaffold hopping and core growing (using the variable-atom coordinates as reference), as well as in scaffold elaboration and linker inpainting (using the full-molecule coordinates to capture the spatial extent of the entire ligand).
The directional covariance encodes a preferred growth direction from a fixed fragment toward a target region. Given the center of mass of the fixed fragment and a target center (the prior center or the center of mass of the variable atoms), the unit direction is
An orthonormal basis is obtained via Gram–Schmidt orthogonalization, yielding a rotation matrix . With an elongation factor (default ) and the trace normalization constraint , the eigenvalues become
and the directional covariance reads
This construction is used for fragment growing, where the direction points from the fixed fragment toward the prior center, and composes with the reference-ligand center-of-mass placement described below.
Both covariance types satisfy the trace invariant , ensuring for , the same second moment as the isotropic prior. This guarantees consistent velocity field magnitudes across prior types and enables stable mixed-batch training in which isotropic and anisotropic samples coexist. Since both and are derived from reference atom coordinates, they transform equivariantly under rigid-body transformations of the protein–ligand complex, preserving the -equivariance of the model. Sampling from is performed via the Cholesky decomposition , yielding with .
Orthogonal to the covariance construction, the prior point cloud can be spatially relocated to improve locality. For global inpainting modes, the prior cloud is shifted to the center of mass of the to-be-generated atoms in the reference ligand. During training, Gaussian noise with standard deviation is added to this placement for robustness. Covariance controls the shape of the prior while placement controls its location; the two compose independently. For substructure and fragment inpainting, the prior cloud is shifted to the center of mass of the to-be-replaced fragment in the reference ligand, and an isotropic prior suffices since the spatial placement itself provides sufficient locality.
Potency and binding affinity prediction
@@
In drug discovery, a variety of parameters, including IC50, Ki, Kd, and EC50, are reported as potency and binding affinity measures. These parameters differ in their definitions and experimental setups, which complicates their direct comparison. For example, Kd is an equilibrium constant that directly measures the strength with which a ligand binds to its target protein. On the other hand, a Ki is an equilibrium constant indicating how well a given inhibitor inhibits the binding of a natural substrate. In other words, while Kd measures the ”stickiness” of a molecule to a target, Ki measures how well the inhibitor blocks the enzyme from its natural substrate. Though these values often correlate, they are not always interchangeable. Likewise, EC50 and IC50 are used for biological assays. Though they correlate with Kd and Ki, their interpretation depends on assay conditions. For better readability, for the remainder of this work we refer to potency and binding affinity prediction as affinity prediction.
@@
Therefore, in contrast to prior work [24, 45], we predict each potency measure separately, aligning better with the complexity of drug-target data. Specifically, Flowr.root employs separate prediction heads for each affinity type (pIC50, pKi, pKd, pEC50), with each head trained separately on its corresponding data type. This design explicitly avoids treating heterogeneous affinity labels as interchangeable. Given invariant and equivariant ligand features , pocket features , and ligand-pocket interaction features extracted from the ligand decoder’s latent message-passing module, the affinity head computes:
@@
where is a Gated equivariant block [67] to combine equivariant and invariant feature tensors, and is a multi-layer perceptron with two linear layers and SiLU activation. The input to is the concatenation of normalized invariant features, where , and denotes the number of atoms in the ligand. The concatenated feature tensor is passed to task-specific heads defined as
@@
The affinity prediction loss is computed using the mean Huber loss between predicted and true affinity values in log units (pIC50, pKd, pKi, pEC50) through
@@
Role of joint structure–affinity training.
We note that the affinity head is architecturally lightweight, comprising four small MLPs that operate on pooled ligand, pocket, and interaction features, totalling 743K parameters (2.3% of the 33M total). The affinity Huber loss constitutes one of eight loss terms; the remaining seven—coordinate MSE, atom-type CE, bond-type CE, charge CE, hybridization CE, bond-length MSE, and bond-angle Huber—collectively dominate the gradient signal to the shared backbone by a substantial margin. While the affinity loss gradients do propagate through the shared representations during training, their relative contribution is minor, consistent with the well-established finding that multi-task learning from related auxiliary objectives provides implicit regularization rather than degradation of the primary task [68, 69]. Multiple lines of empirical evidence underline this: (i) the comparison between Flowr (no affinity head) and Flowr.root (with affinity head), trained from scratch on the same data and protocol, shows that all structural quality improvements (for example, PB-validity, strain energy) are attributable to explicit architectural enhancements (improved equivariant message-passing, residual coordinate updates, bond-length and bond-angle losses) rather than multi-task training; (ii) the unconditional model Flowr.root, which contains no pocket encoder or affinity head, independently achieves leading structural quality on GEOM-Drugs; and (iii) under affinity-guided importance sampling (see Methods), where the affinity head actively steers generation, structural quality metrics remain largely preserved (PB-validity: 0.930.91; strain energy: 48.4654.42 kcal/mol), demonstrating robustness even under maximal affinity-head influence. The purpose of joint training is therefore not to improve structural generation, but to enable a unified workflow supporting in-distribution property-guided generation, joint structure–affinity domain adaptation via LoRA finetuning, multi-target selectivity campaigns, and affinity-annotated fragment elaboration—capabilities that would be lost entirely without the affinity head.
@@
Using distance thresholds and L1 distance errors , the LDDT score for each ligand atom is obtained through
@@
Training and sampling overview
We summarize the training and sampling procedures in Algorithms 1 and 2. The complete ligand representation comprises coordinates , atom types , formal charges , hybridization states , and bond types . For brevity, Algorithms 1–2 detail only coordinates (continuous) and atom types (discrete, categories); all other categorical features follow the identical discrete flow procedure.
Inference-time scaling
Inference-time scaling is a powerful approach that allocates additional computation during generation, rather than solely relying on larger architectures [70, 71]. Originally developed for large language models, this approach has been extended to diffusion and flow matching models. Diffusion models leverage stochasticity through particle sampling [72, 4] and Feynman-Kac steering [73], where multiple trajectories are generated and resampled based on reward functions. Flow matching models, despite being deterministic, achieve inference-time scaling through SDE-based generation, interpolant conversion, and Rollover Budget Forcing (RBF) for adaptive resource allocation [74], improving upon diffusion-based advances [75].
@@
The sequential Monte Carlo (SMC) method [76] extends this concept to a time-sequential setting by maintaining particles and updating their importance weights over time [72]. A straightforward weighting scheme involves evaluating the current batch of particles at time , computing their scores , and obtaining resampling weights through softmax normalization: . Particles are then resampled from the discrete distribution with probability proportional to their weights, as done in Pilot [4].
Quantum Mechanical Calculations on Protein-Ligand Complexes
@@
To benchmark Flowr.root’s affinity prediction head, we perform quantum mechanical (QM) calculations on protein-ligand complexes and correlate the results with predicted affinities. Calculations were conducted using ULYSSES [77], a general-purpose semi-empirical library offering multiple molecular Hamiltonians. We employ GFN2-xTB [78] with ALPB aqueous solvation [79]. GFN2-xTB is a simplified density functional method that parametrizes interparticle interaction integrals against reference data, substantially accelerating calculations [78]—a critical advantage for large atomic systems such as protein-ligand complexes. Additionally, GFN2-xTB provides a balanced treatment of electrostatic interactions and incorporates state-of-the-art dispersion corrections, enabling accurate descriptions of non-bonded complexes. In this work, the QM validation of Flowr.root used full protein domains, containing over 5000 atoms. Such a system size, required for meaningful physics-based evaluations of the granular details distinguishing different ligands, necessitates using semi-empirical Hamiltonians. Further, we only consider non-covalent binders, and, for such cases, intermolecular interactions are treated at a level similar to full Density Functional Theory (DFT). Finally, we found in previous work that GFN2-xTB provides accurate evaluations of protein-ligand complexes, in agreement with experimental SAR data [62, 80, 81, 82].
Regarding the choice of GFN2-xTB for non-covalent protein-ligand binding energetics: (1) The SQM2.20 method demonstrated strong correlations between semi-empirical QM-derived binding energies and experimental data for multiple targets including BACE1, which we also investigate—notably, such correlations were not observed with force-field approaches or contemporary ML scoring functions [85]. (2) GFN2-xTB was explicitly parameterized using non-covalent interaction datasets, employing the D4 dispersion correction and an updated electrostatics model [78], making it particularly suited for non-bonded interactions that dominate protein-ligand binding. (3) Large QM regions are essential for reliably capturing binding energetics and selectivity effects; for example, DYRK1A and DYRK1B are highly homologous kinases (over 85% sequence identity), and treating only small binding-site regions would be insufficient to distinguish relative affinities. We emphasize that the QM calculations serve as an independent validation of predicted affinity trends rather than absolute binding energies.